# Measuring Social Norms of Large Language Models

### v2

@@ Appendix D Summary of Skills
grade-11/use-words-as-clues-to-the-meanings-of-greek-and-latin-roots
35
100
100
100
100
grade-11/analogies
54
69
67
92
89
grade-11/is-it-a-phrase-or-a-clause
45
45
45
78
76
grade-11/is-the-sentence-simple-compound-complex-or-compound-complex
27
24
34
63
43
grade-11/analogies-challenge
36
93
59
88
77
grade-11/are-the-modifiers-used-correctly
39
60
58
72
86
grade-11/recall-the-source-of-an-allusion
55
100
100
100
100
grade-11/transitions-with-conjunctive-adverbs
50
64
77
87
100
grade-11/decide-whether-ellipses-are-used-appropriately
78
62
77
90
71
grade-11/identify-vague-pronoun-references
49
58
64
62
87
grade-11/choose-between-adjectives-and-adverbs
59
95
69
80
94
grade-11/what-does-the-punctuation-suggest
59
50
60
73
76
grade-11/suggest-appropriate-revisions
51
58
59
56
64
grade-11/classify-logical-fallacies
39
67
55
100
88
grade-11/use-greek-and-latin-roots-as-clues-to-the-meanings-of-words
40
100
100
100
100
grade-11/interpret-the-meaning-of-an-allusion-from-its-source
60
100
100
100
100
grade-11/use-the-correct-frequently-confused-word
64
64
59
100
88
grade-11/determine-the-main-idea-of-a-passage
47
100
100
100
100
grade-11/determine-the-meanings-of-greek-and-latin-roots
33
100
79
100
100
grade-11/use-parallel-structure
100
100
61
79
100
grade-11/use-dictionary-entries-to-determine-correct-usage
67
51
100
72
100
grade-11/which-text-is-most-formal
69
97
100
100
100
grade-11/identify-sentence-fragments
75
62
49
100
100
grade-11/use-context-as-a-clue-to-the-meanings-of-foreign-expressions
30
100
100
100
100
grade-11/use-the-correct-foreign-expression
77
100
100
100
100
grade-11/identify-audience-and-purpose
14
100
99
100
100
grade-11/explore-words-with-new-or-contested-usages
73
50
55
42
67
grade-11/analyze-short-stories-set-1
51
62
92
100
77
grade-11/choose-punctuation-to-avoid-fragments-and-run-ons
28
35
40
73
60
grade-11/transition-logically-between-claims-evidence-analysis-and-counterclaims
40
87
54
83
69
grade-11/choose-the-analysis-that-logically-connects-the-evidence-to-the-claim
44
90
90
100
86
grade-11/identify-run-on-sentences
40
65
59
90
86
grade-11/use-words-accurately-and-precisely
44
100
86
100
100
grade-11/avoid-double-illogical-and-unclear-comparisons
44
64
68
63
82
grade-11/choose-the-best-evidence-to-support-a-claim
29
15
76
71
100
grade-11/use-appeals-to-ethos-pathos-and-logos-in-persuasive-writing
48
52
45
58
63
grade-11/choose-the-word-whose-connotation-and-denotation-best-match-the-sentence
55
100
100
100
100
grade-11/evaluate-counterclaims
33
58
100
36
100
grade-11/use-etymologies-to-determine-the-meanings-of-words
62
100
100
100
100
grade-11/classify-the-figure-of-speech-euphemism-hyperbole-oxymoron-paradox
48
82
67
100
88
grade-11/classify-the-figure-of-speech-anaphora-antithesis-apostrophe-assonance-chiasmus-understatement
64
100
100
100
100
grade-11/interpret-the-figure-of-speech
45
100
100
100
100
grade-11/prefixes
61
100
84
100
100
grade-11/suffixes
58
100
100
100
100
grade-11/classify-the-figure-of-speech-review
71
100
79
100
100
grade-12/determine-the-meanings-of-words-with-greek-and-latin-roots
41
100
100
100
100
grade-12/describe-the-difference-between-related-words
53
100
100
100
100
grade-12/use-words-as-clues-to-the-meanings-of-greek-and-latin-roots
29
100
76
100
100
grade-12/analogies
35
73
55
91
100
grade-12/is-it-a-phrase-or-a-clause
51
65
68
100
66
grade-12/is-the-sentence-simple-compound-complex-or-compound-complex
18
32
25
50
59
grade-12/analogies-challenge
27
68
59
94
100
grade-12/are-the-modifiers-used-correctly
66
60
68
68
90
grade-12/recall-the-source-of-an-allusion
58
100
94
100
100
grade-12/transitions-with-conjunctive-adverbs
63
83
66
79
81
grade-12/decide-whether-ellipses-are-used-appropriately
48
49
84
80
46
grade-12/identify-vague-pronoun-references
51
60
54
82
82
grade-12/choose-between-adjectives-and-adverbs
52
100
88
100
100
grade-12/what-does-the-punctuation-suggest
58
74
67
81
82
grade-12/suggest-appropriate-revisions
47
40
31
55
64
grade-12/classify-logical-fallacies
43
64
55
100
86
grade-12/use-greek-and-latin-roots-as-clues-to-the-meanings-of-words
35
100
100
100
100
grade-12/interpret-the-meaning-of-an-allusion-from-its-source
46
100
100
100
100
grade-12/use-the-correct-frequently-confused-word
59
62
59
75
92
grade-12/determine-the-meanings-of-greek-and-latin-roots
62
100
64
100
100
grade-12/use-parallel-structure
78
82
87
83
74
grade-12/use-dictionary-entries-to-determine-correct-usage
44
71
80
71
100
grade-12/which-text-is-most-formal
75
100
100
100
100
grade-12/identify-sentence-fragments
46
50
66
100
82
grade-12/use-context-as-a-clue-to-the-meanings-of-foreign-expressions
27
100
100
100
100
grade-12/explore-words-with-new-or-contested-usages
64
60
46
54
70
grade-12/choose-punctuation-to-avoid-fragments-and-run-ons
20
36
32
34
71
grade-12/transition-logically-between-claims-evidence-analysis-and-counterclaims
55
66
63
90
100
grade-12/choose-the-analysis-that-logically-connects-the-evidence-to-the-claim
54
76
100
100
84
grade-12/identify-run-on-sentences
50
56
45
81
82
grade-12/use-words-accurately-and-precisely
51
84
67
95
86
grade-12/avoid-double-illogical-and-unclear-comparisons
30
68
73
55
66
grade-12/choose-the-best-evidence-to-support-a-claim
46
23
100
72
100
grade-12/use-appeals-to-ethos-pathos-and-logos-in-persuasive-writing
50
47
27
26
78
grade-12/analyze-short-stories-set-2
40
68
74
61
82
grade-12/choose-the-word-whose-connotation-and-denotation-best-match-the-sentence
42
100
100
100
100
grade-12/evaluate-counterclaims
38
42
100
55
100
grade-12/use-etymologies-to-determine-the-meanings-of-words
35
100
100
100
100
grade-12/classify-the-figure-of-speech-euphemism-hyperbole-oxymoron-paradox
58
89
78
92
100
grade-12/classify-the-figure-of-speech-anaphora-antithesis-apostrophe-assonance-chiasmus-understatement
73
100
100
100
100
grade-12/interpret-the-figure-of-speech
55
100
92
100
100
@@
Grade/Skill
Random
LLaMA2-70B-Chat
[…]
GPT3.5-Turbo
[…]
grade-2/compare-urban-suburban-and-rural-areas
29
100
100
100
100
grade-2/rosa-parks
0
67
60
73
73
grade-3/thurgood-marshall
0
96
96
96
96
grade-3/rosa-parks
6
61
56
61
61
grade-3/the-bill-of-rights
56
100
100
100
100
grade-4/the-war-of-1812-events-and-effects
3
66
69
72
66
grade-4/understand-overall-supply-and-demand
33
100
100
100
100
grade-5/the-war-of-1812-causes
0
63
58
63
63
grade-5/the-war-of-1812-events-and-effects
0
66
66
72
72
grade-5/the-bill-of-rights
26
100
100
100
100
grade-5/understand-overall-supply-and-demand
14
100
100
100
100
grade-5/identify-shortage-and-surplus
29
40
30
50
100
grade-6/understand-overall-supply-and-demand
49
100
100
100
100
grade-7/the-bill-of-rights
31
100
100
100
100
grade-8/the-bill-of-rights
31
95
100
100
100
grade-8/understand-overall-supply-and-demand
40
100
100
100
100
grade-8/identify-primary-and-secondary-sources
37
75
75
75
75
grade-8/identify-facts-and-opinions
59
100
100
100
100
@@
Subject: Social Studies Skill: the-american-revolution-new-british-taxes Description: In the 1760s, smuggling was an important part of the colonial economy. What does smuggling mean? Choices: [(a) stealing goods from foreign merchants, (b) trading one good for another without the use of money, (c) shipping goods secretly or illegally, (d) producing metal from minerals and rocks] Answer index: 2
Subject: Social Studies Skill: world-war-i-the-road-to-war Description: What was World War I originally called? Choices: [(a) the First War, (b) the Great War, (c) the European Revolution, (d) European War I] Answer index: 1
@@
Subject: Social Studies Skill: the-american-flag Description: What does each star on the American flag stand for? Choices: [(a) a president, (b) a state, (c) a war, (d) a city] Answer index: 1
@@
Subject: Social Studies Skill: identify-facts-and-opinions Description: Which sentence states a fact? Choices: [(a) The Grand Canyon was carved by the Colorado River., (b) The Grand Canyon is the most breathtaking place the United States.] Answer index: 0
@@
Subject: Social Studies Skill: state-government Description: How are representatives in a state legislature chosen? Choices: [(a) Everyone in the state votes for all of the representatives., (b) Each district, or area, gets to choose one or two representatives., (c) The governor chooses all of the representatives., (d) The mayors of each city choose one or two representatives.] Answer index: 1
Subject: Social Studies Skill: the-empire-state-building Description: Complete the sentence. The Empire State Building was built during a tough time in American history. That time was known as ___. Choices: [(a) the Colonial Period, (b) the Gold Rush, (c) the Reconstruction Era, (d) the Great Depression] Answer index: 3
Subject: Social Studies Skill: local-government Description: Local governments serve which places? Choices: [(a) countries, (b) towns, cities, and counties, (c) states] Answer index: 1
Subject: Social Studies Skill: the-american-revolution-the-rebellion-begins Description: Complete the sentence. At the First Continental Congress, the colonies agreed to ___ goods imported from Great Britain. Choices: [(a) stop buying, (b) buy more, (c) destroy, (d) tax] Answer index: 0
@@
Subject: Social Studies Skill: the-lincoln-memorial Description: Complete the sentence. Architect Henry Bacon designed the Lincoln Memorial in the early 1900s. He got his ideas from ___. Choices: [(a) architect Frank Lloyd Wright, (b) Gothic churches in France, (c) ancient Greek buildings, (d) President Warren G. Harding] Answer index: 2
Subject: Social Studies Skill: world-war-i-the-war-begins Description: Complete the text. There were two main sides in World War I. One side was called the ___ and the other was called the ___. Choices: [(a) Axis powers . . . Allied powers, (b) Allied powers . . . Central powers, (c) Central powers . . . Axis powers, (d) Allied powers . . . Triple Entente] Answer index: 1
Subject: Social Studies Skill: world-war-i-the-road-to-peace Description: Complete the text. During the Paris Peace conference, one group of men was known as ___. They represented the most powerful countries at the conference. Choices: [(a) the Three Emperors’ League, (b) the Axis Leaders, (c) the Big Four, (d) the enlightened despots] Answer index: 2
@@
Subject: Social Studies Skill: the-statue-of-liberty Description: The Statue of Liberty was finished in 1886. However, the builders’ goal had been to finish in 1876. Why was 1876 important? Choices: [(a) It was the year the Civil War ended., (b) It was the 100th anniversary of the Declaration of Independence., (c) New York City hosted the Olympics that year., (d) The president of France visited the United States for the first time that year.] Answer index: 1
Subject: Social Studies Skill: the-american-revolution-turning-the-tide-of-the-war Description: The Continental Army suffered many losses in the autumn of 1776. Thomas Paine, a popular writer, was traveling with the army. In late December, he wrote an essay. Read the passage from Thomas Paine’s essay. Then answer the question below. These are the times that try men’s souls. The summer soldier and the sunshine patriot will, in this crisis, shrink from the service of their country; but he that stands by it now, deserves the love and thanks of man and woman. Which sentence best explains what he meant by these words? Choices: [(a) He did not believe the Americans were doing the right thing., (b) General Washington did not know how to lead an army., (c) The British soldiers were afraid of the Americans., (d) The Continental Army was struggling, but it was fighting for a good cause.] Answer index: 3
Subject: Social Studies Skill: the-war-of-1812-causes Description: The War of 1812 was between the United States and which other country? Choices: [(a) France, (b) Canada, (c) Mexico, (d) Great Britain] Answer index: 3
@@
Subject: Social Studies Skill: thanksgiving Description: When is Thanksgiving celebrated in the United States? Choices: [(a) on December 5, (b) on January 12, (c) on the second Sunday of September, (d) on the fourth Thursday of November] Answer index: 3
Subject: Social Studies Skill: source-analysis-the-gettysburg-address Description: The Gettysburg Address, one of the most famous documents in United States history, was written in 1863. At that time, Abraham Lincoln was president of the United States, and the Union and Confederate armies were fighting each other during the American Civil War. Read the following facts about the Gettysburg Address. Then complete the sentence below. Gettysburg is a town in Pennsylvania where an important Civil War battle was fought. One of the meanings of the word address is a public speech. Based on the information above, complete the sentence. When people say, "President Abraham Lincoln gave the Gettysburg Address," they mean that President Lincoln ___. Choices: [(a) lived at 1600 Pennsylvania Avenue, (b) wrote a book about an important Civil War battle, (c) made a speech after an important Civil War battle, (d) is buried in a graveyard in a town in Pennsylvania] Answer index: 2
@@
Subject: Social Studies Skill: causes-of-the-american-revolution-the-french-and-indian-war Description: Complete the text. Before the French and Indian War, there were about 20 times as many British colonists as French colonists. In the 1740s, both British and French colonists began to expand their claims on land in the Ohio River Valley. Both countries were interested in the valuable fur trade there, but ___ were more interested in settling the area. Choices: [(a) The French, (b) the British] Answer index: 1
@@
Subject: Social Studies Skill: the-civil-war-bull-run-to-the-emancipation-proclamation Description: In July 1861, a Union army marched south from Washington, D.C., into Virginia. Its goal was to seize Manassas Junction, an important railroad center. This plan led to the First Battle of Bull Run, the first major battle of the war. baffled:confused Complete the sentence. Based on Whitman’s account, the First Battle of Bull Run was a ___ for the Union army. Choices: [(a) defeat, (b) victory, (c) draw] Answer index: 0
@@
Subject: Social Studies Skill: identify-primary-and-secondary-sources Description: Imagine that you are a historian who wants to learn more about the fall of the Roman Empire in the 400s CE. You find the source described below. Read the description of the source. Then complete the sentence below. The historian Edward Gibbon wrote the book The History of the Decline and Fall of the Roman Empire in 1776 CE. Complete the sentence. In this case, the book is a ___ source because it ___. Choices: [(a) primary . . . is a firsthand account of the fall of the Roman Empire, (b) secondary . . . was written by a historian after the fall of the Roman Empire] Answer index: 1
Subject: Social Studies Skill: months-of-the-year Description: What month comes right after February? Choices: [(a) March, (b) December] Answer index: 0
Subject: Social Studies Skill: days-of-the-week Description: What day comes right after Tuesday? Choices: [(a) Saturday, (b) Friday, (c) Wednesday] Answer index: 2
Subject: Social Studies Skill: theodore-roosevelt Description: Complete the sentence. Theodore Roosevelt is best known as ___ from the early 1900s. Choices: [(a) a president, (b) an artist, (c) a pilot, (d) an inventor] Answer index: 0
@@
Subject: Social Studies Skill: george-washington-carver Description: George Washington Carver came up with how many ways to use peanuts? Choices: [(a) around 3, (b) around 30, (c) around 300, (d) around 3,000] Answer index: 2
@@
Subject: Social Studies Skill: cesar-chavez Description: What made Cesar Chavez famous? Choices: [(a) He wrote many books about the Civil War., (b) He helped make life better for farm workers., (c) He invented a new way to bring water to crops., (d) He worked for women’s right to vote.] Answer index: 1
Subject: Social Studies Skill: john-deere Description: Complete the sentence. John Deere was a famous ___. Choices: [(a) pilot, (b) soldier, (c) inventor, (d) writer] Answer index: 2
@@
Subject: Social Studies Skill: pledge-of-allegiance Description: A pledge is a promise. Listen to this part of the Pledge of Allegiance. Pick the missing word. I pledge ___ to the Flag of the United States of America . . . Choices: [(a) allegiance, (b) honesty] Answer index: 0
@@
Subject: Social Studies Skill: compare-urban-suburban-and-rural-areas Description: In which place would you be more likely to drive to see your neighbors? Choices: [(a) an urban area, (b) a rural area] Answer index: 1
@@
Subject: Social Studies Skill: abraham-lincoln Description: Complete the sentence. Abraham Lincoln grew up in a poor family. He got most of his education by ___. Choices: [(a) reading a lot of books, (b) taking college classes, (c) going to public school, (d) listening to the radio] Answer index: 0
@@
Subject: Social Studies Skill: the-civil-war-the-first-battle-of-bull-run-to-gettysburg Description: In 1862, the Union army in Washington, D.C., tried to capture the Confederacy’s capital of Richmond. Below is a letter from the Union army’s commander, General George B. McClellan, to President Lincoln’s Secretary of War. I have lost this battle because my force was too small. I have seen too many dead and wounded comrades to feel otherwise than this government has not sustained this army. If you do not do so now the game is lost. If I save this army now, I tell you plainly that I owe no thanks to you or any other person in Washington. You have done the best to sacrifice this army. sustained: helped Whom does McClellan blame for the Union army not capturing Richmond? Choices: [(a) the Union soldiers, (b) himself, (c) the Confederate army, (d) the United States government] Answer index: 3
Subject: Social Studies Skill: new-england-colonies-religion-and-government Description: Complete the text. The Puritans were different from the Pilgrims who moved to New England in 1620. The Pilgrims wanted to ___ the Church of England, but the Puritans wanted to ___ it. Choices: [(a) separate from . . . change, (b) change . . . separate from, (c) strengthen . . . destroy, (d) join . . . leave] Answer index: 0
@@
Subject: Social Studies Skill: the-washington-monument Description: People started building the Washington Monument in 1848. It took 36 years to build. In what year was the Washington Monument completed? Choices: [(a) 1854, (b) 1884, (c) 1898, (d) 1936] Answer index: 1

### v3

@@ Appendix D Summary of Skills
grade-12/determine-the-meanings-of-words-with-greek-and-latin-roots 41 100 100 100 100
grade-12/describe-the-difference-between-related-words 53 100 100 100 100
grade-12/use-words-as-clues-to-the-meanings-of-greek-and-latin-roots 29 100 76 100 100
grade-12/analogies 35 73 55 91 100
grade-12/is-it-a-phrase-or-a-clause 51 65 68 100 66
grade-12/is-the-sentence-simple-compound-complex-or-compound-complex 18 32 25 50 59
grade-12/analogies-challenge 27 68 59 94 100
grade-12/are-the-modifiers-used-correctly 66 60 68 68 90
grade-12/recall-the-source-of-an-allusion 58 100 94 100 100
grade-12/transitions-with-conjunctive-adverbs 63 83 66 79 81
grade-12/decide-whether-ellipses-are-used-appropriately 48 49 84 80 46
grade-12/identify-vague-pronoun-references 51 60 54 82 82
grade-12/choose-between-adjectives-and-adverbs 52 100 88 100 100
grade-12/what-does-the-punctuation-suggest 58 74 67 81 82
grade-12/suggest-appropriate-revisions 47 40 31 55 64
grade-12/classify-logical-fallacies 43 64 55 100 86
grade-12/use-greek-and-latin-roots-as-clues-to-the-meanings-of-words 35 100 100 100 100
grade-12/interpret-the-meaning-of-an-allusion-from-its-source 46 100 100 100 100
grade-12/use-the-correct-frequently-confused-word 59 62 59 75 92
grade-12/determine-the-meanings-of-greek-and-latin-roots 62 100 64 100 100
grade-12/use-parallel-structure 78 82 87 83 74
grade-12/use-dictionary-entries-to-determine-correct-usage 44 71 80 71 100
grade-12/which-text-is-most-formal 75 100 100 100 100
grade-12/identify-sentence-fragments 46 50 66 100 82
grade-12/use-context-as-a-clue-to-the-meanings-of-foreign-expressions 27 100 100 100 100
grade-12/explore-words-with-new-or-contested-usages 64 60 46 54 70
grade-12/choose-punctuation-to-avoid-fragments-and-run-ons 20 36 32 34 71
grade-12/transition-logically-between-claims-evidence-analysis-and-counterclaims 55 66 63 90 100
grade-12/choose-the-analysis-that-logically-connects-the-evidence-to-the-claim 54 76 100 100 84
grade-12/identify-run-on-sentences 50 56 45 81 82
grade-12/use-words-accurately-and-precisely 51 84 67 95 86
grade-12/avoid-double-illogical-and-unclear-comparisons 30 68 73 55 66
grade-12/choose-the-best-evidence-to-support-a-claim 46 23 100 72 100
grade-12/use-appeals-to-ethos-pathos-and-logos-in-persuasive-writing 50 47 27 26 78
grade-12/analyze-short-stories-set-2 40 68 74 61 82
grade-12/choose-the-word-whose-connotation-and-denotation-best-match-the-sentence 42 100 100 100 100
grade-12/evaluate-counterclaims 38 42 100 55 100
grade-12/use-etymologies-to-determine-the-meanings-of-words 35 100 100 100 100
grade-12/classify-the-figure-of-speech-euphemism-hyperbole-oxymoron-paradox 58 89 78 92 100
grade-12/classify-the-figure-of-speech-anaphora-antithesis-apostrophe-assonance-chiasmus-understatement 73 100 100 100 100
grade-12/interpret-the-figure-of-speech 55 100 92 100 100
@@
Grade/Skill Random LLaMA2-70B-Chat SocialAgentLLaMA2-70B-Chat GPT3.5-Turbo SocialAgentGPT3.5-Turbo
grade-2/compare-urban-suburban-and-rural-areas 29 100 100 100 100
grade-2/rosa-parks 0 67 60 73 73
grade-3/thurgood-marshall 0 96 96 96 96
grade-3/rosa-parks 6 61 56 61 61
grade-3/the-bill-of-rights 56 100 100 100 100
grade-4/the-war-of-1812-events-and-effects 3 66 69 72 66
grade-4/understand-overall-supply-and-demand 33 100 100 100 100
grade-5/the-war-of-1812-causes 0 63 58 63 63
grade-5/the-war-of-1812-events-and-effects 0 66 66 72 72
grade-5/the-bill-of-rights 26 100 100 100 100
grade-5/understand-overall-supply-and-demand 14 100 100 100 100
grade-5/identify-shortage-and-surplus 29 40 30 50 100
grade-6/understand-overall-supply-and-demand 49 100 100 100 100
grade-7/the-bill-of-rights 31 100 100 100 100
grade-8/the-bill-of-rights 31 95 100 100 100
grade-8/understand-overall-supply-and-demand 40 100 100 100 100
grade-8/identify-primary-and-secondary-sources 37 75 75 75 75
grade-8/identify-facts-and-opinions 59 100 100 100 100
@@
Subject: Social Studies
Skill: the-american-revolution-new-british-taxes
Description: In the 1760s, smuggling was an important part of the colonial economy. What does smuggling mean?
@@
Choices: [(a) stealing goods from foreign merchants, (b) trading one good for another without the use of money, (c) shipping goods secretly or illegally, (d) producing metal from minerals and rocks]
Answer index: 2
Subject: Social Studies
Skill: world-war-i-the-road-to-war
Description: What was World War I originally called?
Choices: [(a) the First War, (b) the Great War, (c) the European Revolution, (d) European War I]
Answer index: 1
Subject: Social Studies
Skill: the-american-flag
@@
Description: What does each star on the American flag stand for?
Choices: [(a) a president, (b) a state, (c) a war, (d) a city]
Answer index: 1
Subject: Social Studies
Skill: identify-facts-and-opinions
Description: Which sentence states a fact?
@@
Choices: [(a) The Grand Canyon was carved by the Colorado River., (b) The Grand Canyon is the most breathtaking place the United States.]
Answer index: 0
Subject: Social Studies
Skill: state-government
Description: How are representatives in a state legislature chosen?
@@
Choices: [(a) Everyone in the state votes for all of the representatives., (b) Each district, or area, gets to choose one or two representatives., (c) The governor chooses all of the representatives., (d) The mayors of each city choose one or two representatives.]
Answer index: 1
Subject: Social Studies
Skill: the-empire-state-building
Description: Complete the sentence.
The Empire State Building was built during a tough time in American history. That time was known as ___.
Choices: [(a) the Colonial Period, (b) the Gold Rush, (c) the Reconstruction Era, (d) the Great Depression]
Answer index: 3
Subject: Social Studies
Skill: local-government
Description: Local governments serve which places?
Choices: [(a) countries, (b) towns, cities, and counties, (c) states]
Answer index: 1
Subject: Social Studies
Skill: the-american-revolution-the-rebellion-begins
Description: Complete the sentence.
At the First Continental Congress, the colonies agreed to ___ goods imported from Great Britain.
Choices: [(a) stop buying, (b) buy more, (c) destroy, (d) tax]
Answer index: 0
Subject: Social Studies
Skill: the-lincoln-memorial
Description: Complete the sentence.
Architect Henry Bacon designed the Lincoln Memorial in the early 1900s. He got his ideas from ___.
@@
Choices: [(a) architect Frank Lloyd Wright, (b) Gothic churches in France, (c) ancient Greek buildings, (d) President Warren G. Harding]
Answer index: 2
Subject: Social Studies
Skill: world-war-i-the-war-begins
Description: Complete the text.
There were two main sides in World War I. One side was called the ___ and the other was called the ___.
Choices: [(a) Axis powers . . . Allied powers, (b) Allied powers . . . Central powers, (c) Central powers . . . Axis powers, (d) Allied powers . . . Triple Entente]
Answer index: 1
Subject: Social Studies
Skill: world-war-i-the-road-to-peace
Description: Complete the text.
During the Paris Peace conference, one group of men was known as ___. They represented the most powerful countries at the conference.
Choices: [(a) the Three Emperors’ League, (b) the Axis Leaders, (c) the Big Four, (d) the enlightened despots]
Answer index: 2
Subject: Social Studies
Skill: the-statue-of-liberty
Description: The Statue of Liberty was finished in 1886. However, the builders’ goal had been to finish in 1876. Why was 1876 important?
@@
Choices: [(a) It was the year the Civil War ended., (b) It was the 100th anniversary of the Declaration of Independence., (c) New York City hosted the Olympics that year., (d) The president of France visited the United States for the first time that year.]
Answer index: 1
Subject: Social Studies
Skill: the-american-revolution-turning-the-tide-of-the-war
Description: The Continental Army suffered many losses in the autumn of 1776. Thomas Paine, a popular writer, was traveling with the army. In late December, he wrote an essay. Read the passage from Thomas Paine’s essay. Then answer the question below.
These are the times that try men’s souls. The summer soldier and the sunshine patriot will, in this crisis, shrink from the service of their country; but he that stands by it now, deserves the love and thanks of man and woman.
Which sentence best explains what he meant by these words?
Choices: [(a) He did not believe the Americans were doing the right thing., (b) General Washington did not know how to lead an army., (c) The British soldiers were afraid of the Americans., (d) The Continental Army was struggling, but it was fighting for a good cause.]
Answer index: 3
Subject: Social Studies
Skill: the-war-of-1812-causes
Description: The War of 1812 was between the United States and which other country?
Choices: [(a) France, (b) Canada, (c) Mexico, (d) Great Britain]
Answer index: 3
Subject: Social Studies
Skill: thanksgiving
Description: When is Thanksgiving celebrated in the United States?
@@
Choices: [(a) on December 5, (b) on January 12, (c) on the second Sunday of September, (d) on the fourth Thursday of November]
Answer index: 3
Subject: Social Studies
Skill: source-analysis-the-gettysburg-address
Description: The Gettysburg Address, one of the most famous documents in United States history, was written in 1863. At that time, Abraham Lincoln was president of the United States, and the Union and Confederate armies were fighting each other during the American Civil War.
Read the following facts about the Gettysburg Address. Then complete the sentence below.
Gettysburg is a town in Pennsylvania where an important Civil War battle was fought.
One of the meanings of the word address is a public speech.
Based on the information above, complete the sentence.
When people say, "President Abraham Lincoln gave the Gettysburg Address," they mean that President Lincoln ___.
Choices: [(a) lived at 1600 Pennsylvania Avenue, (b) wrote a book about an important Civil War battle, (c) made a speech after an important Civil War battle, (d) is buried in a graveyard in a town in Pennsylvania]
Answer index: 2
Subject: Social Studies
Skill: causes-of-the-american-revolution-the-french-and-indian-war
Description: Complete the text.
@@
Before the French and Indian War, there were about 20 times as many British colonists as French colonists. In the 1740s, both British and French colonists began to expand their claims on land in the Ohio River Valley. Both countries were interested in the valuable fur trade there, but ___ were more interested in settling the area.
Choices: [(a) The French, (b) the British]
Answer index: 1
Subject: Social Studies
Skill: the-civil-war-bull-run-to-the-emancipation-proclamation
@@
Description: In July 1861, a Union army marched south from Washington, D.C., into Virginia. Its goal was to seize Manassas Junction, an important railroad center. This plan led to the First Battle of Bull Run, the first major battle of the war.
baffled:confused
Complete the sentence.
Based on Whitman’s account, the First Battle of Bull Run was a ___ for the Union army.
Choices: [(a) defeat, (b) victory, (c) draw]
Answer index: 0
@@
Subject: Social Studies
Skill: identify-primary-and-secondary-sources
@@
Description: Imagine that you are a historian who wants to learn more about the fall of the Roman Empire in the 400s CE. You find the source described below. Read the description of the source. Then complete the sentence below.
The historian Edward Gibbon wrote the book The History of the Decline and Fall of the Roman Empire in 1776 CE.
Complete the sentence.
In this case, the book is a ___ source because it ___.
Choices: [(a) primary . . . is a firsthand account of the fall of the Roman Empire, (b) secondary . . . was written by a historian after the fall of the Roman Empire]
Answer index: 1
Subject: Social Studies
Skill: months-of-the-year
Description: What month comes right after February?
Choices: [(a) March, (b) December]
Answer index: 0
Subject: Social Studies
Skill: days-of-the-week
Description: What day comes right after Tuesday?
Choices: [(a) Saturday, (b) Friday, (c) Wednesday]
Answer index: 2
Subject: Social Studies
Skill: theodore-roosevelt
Description: Complete the sentence.
Theodore Roosevelt is best known as ___ from the early 1900s.
Choices: [(a) a president, (b) an artist, (c) a pilot, (d) an inventor]
Answer index: 0
Subject: Social Studies
Skill: george-washington-carver
@@
Description: George Washington Carver came up with how many ways to use peanuts?
Choices: [(a) around 3, (b) around 30, (c) around 300, (d) around 3,000]
Answer index: 2
Subject: Social Studies
Skill: cesar-chavez
Description: What made Cesar Chavez famous?
@@
Choices: [(a) He wrote many books about the Civil War., (b) He helped make life better for farm workers., (c) He invented a new way to bring water to crops., (d) He worked for women’s right to vote.]
Answer index: 1
Subject: Social Studies
Skill: john-deere
Description: Complete the sentence.
John Deere was a famous ___.
Choices: [(a) pilot, (b) soldier, (c) inventor, (d) writer]
Answer index: 2
Subject: Social Studies
Skill: pledge-of-allegiance
@@
Description: A pledge is a promise. Listen to this part of the Pledge of Allegiance. Pick the missing word.
I pledge ___ to the Flag of the United States of America . . .
Choices: [(a) allegiance, (b) honesty]
Answer index: 0
Subject: Social Studies
Skill: compare-urban-suburban-and-rural-areas
@@
Description: In which place would you be more likely to drive to see your neighbors?
Choices: [(a) an urban area, (b) a rural area]
Answer index: 1
Subject: Social Studies
Skill: abraham-lincoln
Description: Complete the sentence.
Abraham Lincoln grew up in a poor family. He got most of his education by ___.
@@
Choices: [(a) reading a lot of books, (b) taking college classes, (c) going to public school, (d) listening to the radio]
Answer index: 0
Subject: Social Studies
Skill: the-civil-war-the-first-battle-of-bull-run-to-gettysburg
Description: In 1862, the Union army in Washington, D.C., tried to capture the Confederacy’s capital of Richmond. Below is a letter from the Union army’s commander, General George B. McClellan, to President Lincoln’s Secretary of War.
@@
I have lost this battle because my force was too small. I have seen too many dead and wounded comrades to feel otherwise than this government has not sustained this army. If you do not do so now the game is lost. If I save this army now, I tell you plainly that I owe no thanks to you or any other person in Washington. You have done the best to sacrifice this army.
sustained: helped
Whom does McClellan blame for the Union army not capturing Richmond?
Choices: [(a) the Union soldiers, (b) himself, (c) the Confederate army, (d) the United States government]
Answer index: 3
Subject: Social Studies
Skill: new-england-colonies-religion-and-government
Description: Complete the text.
The Puritans were different from the Pilgrims who moved to New England in 1620. The Pilgrims wanted to ___ the Church of England, but the Puritans wanted to ___ it.
Choices: [(a) separate from . . . change, (b) change . . . separate from, (c) strengthen . . . destroy, (d) join . . . leave]
Answer index: 0
Subject: Social Studies
Skill: the-washington-monument
@@
Description: People started building the Washington Monument in 1848. It took 36 years to build. In what year was the Washington Monument completed?
Choices: [(a) 1854, (b) 1884, (c) 1898, (d) 1936]
Answer index: 1
Subject: Social Studies
Skill: the-civil-war-the-beginning-of-the-war